OpenStudy (anonymous):

7/8 p − 4 = 10 for p

OpenStudy (anonymous):

7/8p=14

OpenStudy (anonymous):

14 * 8/7

OpenStudy (anonymous):

112/7

OpenStudy (anonymous):

= ?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

7/8-4=10 +4 +4 ____________________ (8/7) 7/8p=10+4(8/7) p= 16

OpenStudy (anonymous):

add the four over, then you get 7/8p=10+4 the mult the reciprocal of 7/8 on both sides, which is 8/7....the with the variable cancels then continue working to find p and 8/7 * 10+4=16

OpenStudy (anonymous):

7/8p-4/1=10 7p-32/8=10 7p-32=80 7p=80+32 7p=112 P=112/7 P=16

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

NP

OpenStudy (anonymous):

:-)